Part B

Question 8 In the middle of the night you are stan
Answer #1

horizontal distance x= horizontal velocity * time taken


x = 13.2*cos(54)*T


T is the time of flight


using the equation s = (uy*t)- (0.5*g*T^2)

-1.6 = (13.2*sin(54)*T)-(0.5*9.81*T^2)

T = 2.32 s


then x = 13.2*cos(54)*2.32 =    18 m


So the required distamce is d = 18-14 = 4 m